It takes 5 men to paint 2 identical houses in 3 days. Assuming that all men work at the same rate, how long will it take 10 men to paint 8 of these houses?
Answer by stanbon(75887)

It takes 5 men to paint 2 identical houses in 3 days. Assuming that all men work at the same rate, how long will it take 10 men to paint 8 of these houses?
----
time is directly related to # of houses and indirectly related to # of men
-----
t = k*h/m
-----------
Solve for "k" using "it takes 5 men to paint 2 houses in 3 days".
3 = k*2/5
k = 15/2 = 7.5
--------
Equation:
t = 7.5*h/m
----
how long will it take 10 men to paint 8 of these houses?
time = 7.5*8/10
----
time = 6 days
===================
Cheers,
Stan H.
================---